# A Reduction-Driven Local Search for the Generalized Independent Set Problem

### v1

@@ A Technical Proofs
When $w(u)\geq p(u,v)$ and $w(u)\geq 0$. If $u\notin I$ and $v\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)-p(u,v)\geq nb(I)$. If $u\notin I$ and $v\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$. In summary, $u$ is included in some set in $MGIS(G)$.
When $w(u)<0$ and $w(u)\geq p(u,v)$. If $v\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)-p(u,v)\geq nb(I)$ for $u\notin I$. If $v\notin I$, then $nb(I\cup\{u\})<nb(I\setminus\{u\})$. In other word, there exists a set $S\in MGIS(G)$ that $u\in S$ if and only if there exists a set $S^{\prime}\in MGIS(G^{\prime})$ that $v\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,v$ that could appear in a set in $MGIS(G)$.
When $w(u)\geq 0$, $w(u)<p(u,v)$ and $w(u)\geq\tilde{w}(v)$. If $u\notin I,v\in I$, then $nb(I\cup\{u\})\setminus\{v\})-nb(I)=w(u)-w(v)-\sum_{x\in N_{r}(v)\cap I}max(0,-p(v,x))=w(u)-\tilde{w}(v)\geq 0$, thus $I\cup\{u\}\in MGIS(G)$. Otherwise, if $u\notin I,v\notin I$, then $nb(I\cup\{u\})-nb(I)=w(u)\geq 0$, thus $I\cup\{u\}\in MGIS(G)$. In summary, $u$ is included in some vertex set $I^{\prime}\in MGIS(G)$ and $v$ is not in $I^{\prime}$.
When $w(u)\geq 0$, $w(u)<p(u,v)$ and $w(u)<\tilde{w}(v)$. $u$ and $v$ cannot be both included in $I$. Because if $u,v\in I$, then $nb(I\setminus\{u\})=nb(I)+p(u,v)-w(u)>nb(I)$ contradicts to the optimality of $I$. Furthermore, if $v\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I\setminus\{u\})=w(u)\geq 0$. We conclude that there exists a vertex set in $MGIS(G)$ that includes $u$ iff there exists a vertex set in $MGIS(G^{\prime})$ that does not contain $v$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,v$ that could appear in a set in $MGIS(G)$.
Finally, when $w(u)<0$ and $w(u)<p(u,v)$. We have $nb(I\cup\{u\})<nb(I\setminus\{u\})$, therefore $u$ is not in any set in $MGIS(G)$ and we can directly remove $u$. ∎
@@
Given an arbitrary $I\in MGIS(G)$. Because that $p(x,y)=M$, it holds that $x$ and $y$ cannot be both included in $I$.
Case 1: $w(u)\geq\max(p(u,x),p(u,y))$.
When $w(u)\geq 0$. If both $x$ and $y$ are not included in $I$ and $u\notin I$, then $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$. If $u\notin I$ and either $x\in I$ or $y\in I$, $nb(I\cup\{u\})\geq nb(I)+w(u)-\max(p(u,x),p(u,y))\geq nb(I)$. In summary, there exists a vertex set in $MGIS(G)$ that includes $u$.
When $w(u)<0$. If both $x$ and $y$ are not included in $I$, then $u\notin I$ because $nb(I\cup\{u\})=nb(I\setminus\{u\})+w(u)<nb(I\setminus\{u\})$. If either $x\in I$ or $y\in I$ and $u\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})\geq nb(I)+w(u)-\max(p(u,x),p(u,y))\geq nb(I)$. We conclude that there exists a vertex set in $MGIS(G)$ that includes $u$ iff there exists a vertex set in $MGIS(G^{\prime})$ that includes either $x$ or $y$. In other word, we can defer the decision of $u$ after determining $x$ and $y$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
Case 2: $p(u,y)\leq w(u)<p(u,x)$. In this case, if $y$ is in $I$, then $x\notin I$ and $I\cup\{u\}\in MGIS(G)$.
@@
When $w(u)\geq 0$ and $w(u)<\tilde{w}(x)$. There are five cases: $x\in I$ and $u,y\notin I$, $y\in I$ and $u,x\notin I$, $u\in I$ and $x,y\notin I$, $u,y\in I$ and $x\notin I$, and $u,x\in I$ and $y\notin I$. But $u,x$ cannot appear in same $I$, because otherwise $nb(I\setminus\{u\})=nb(I)+p(u,x)-w(u)>0$ contradicts the optimality of $I$. And when $y\in I$ and $u,x\notin I$, then $I\cup\{u\}\in MGIS(G)$. In summary, there exists a vertex set $I^{\prime}\in MGIS(G)$ such that either $x\in I^{\prime}$, or both $u,y\in I^{\prime}$, or $u\in I^{\prime}$. Since $u$ only connects to $x$ and $y$ in $G$, we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include $x$. In other word, we can defer $u$ after the decision of $x$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
When $w(u)<0$. If $y\notin I$, then $u\notin I$ because otherwise $nb(I\setminus\{u\})>nb(I)$. Since $u$ only connects to $x$ and $y$ in $G$, we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that includes $y$. In other word, we can defer $u$ after the decision of $y$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
Case 3: $w(u)<p(u,x)$ and $w(u)<p(u,y)$. When $w(u)\geq 0$, the three vertices form a triangle with permanent edges. The result follows from theorem 3.3 of Gu et al. (2021). In particular, (1) if $w(u)\geq 0$ and $w(u)\geq\tilde{w}(x)$, then $u$ must be in some set in $MGIS(G)$. (2) If $w(u)\geq 0$ and $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$, then $y$ must not be in some set in $MGIS(G)$. And there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include $x$. (3) If $w(u)\geq 0$ and $w(u)<\tilde{w}(y)$, then there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include both $x$ and $y$. (4) When $w(u)<0$, the result follows from R 5. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all conditions and all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$. ∎
@@
Case 1: $w(u)\geq p(u,x)$. In this case, if one of $x,y$ is in $I$, say $x$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I)\geq w(u)-p(u,x)\geq 0$ when $u\notin I$.
@@
When $w(u)\geq 0$ and $w(u)\geq p(u,x)+p(u,y)$. If either $x\in I$ or $y\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I)\geq w(u)-\max(p(u,x)+p(u,y),p(u,x))\geq 0$ for $u\notin I$. If none of $x,y$ is in $I$, then $I\cup\{u\}\in MGIS(G)$ because $w(u)\geq 0$. Therefore, $u$ must belong to some set in $MGIS$.
@@
When $w(u)\geq 0$ and $w(u)<p(u,x)+p(u,y)$. If both $x,y$ are in $I$, then $nb(I\cup\{u\})<nb(I)$. If either $x\notin I$ or $y\notin I$, then $nb(I\cup\{u\})\geq nb(I)$. In other word, including $u$ has negative profit only if both $x$ and $y$ are included in the $MGIS$. we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that either $x\notin S^{\prime}$ or $y\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
@@
When $w(u)<0$, it holds that $w(u)\geq p(u,x)+p(u,y)$ because $w(u)\geq p(u,x)$ and $p(u,y)<0$. If either $x\in I$ or $y\in I$, then $nb(I\cup\{u\})\geq nb(I)$. If both $x\notin I$ and $y\notin I$, then $nb(I\cup\{u\})<nb(I)$. In other word, including $u$ has positive profit if and only if either $x$ or $y$ is included. we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that either $x\in S^{\prime}$ or $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
Case 2: $p(u,y)\leq w(u)<p(u,x)$. In this case, we consider whether the three vertices $u,x,y$ can be included in $I$.
@@
When $w(u)\geq 0$ and $w(u)\geq p(u,x)+p(u,y)$, the all three vertices may be included in $I$. All combinations of $x,y,u$ may appear in $I$. Except that if $x$ is included and $y$ is not included in $I$, then $u\notin I$ because $nb(I\cup\{u\})=nb(I\setminus\{u\})+w(u)-p(u,x)<nb(I\setminus\{u\})$. In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that either $x\notin S^{\prime}$ or $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
@@
When $w(u)\geq 0$ and $w(u)<p(u,x)+p(u,y)$, the three vertices can not be included in $I$. Moreover, $u$ and $x$ cannot be both included by $I$ because otherwise $nb(I\setminus u)>nb(I)$ contradicts the optimality of $I$. If $x\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})\geq nb(I\setminus\{u\})$. In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
@@
When $w(u)<0$ and $w(u)\geq p(u,x)+p(u,y)$, the three vertices can be included in $I$. If $y\notin I$, then $nb(I\cup\{u\})\leq nb(I\setminus\{u\})$ regardless of whether $x\in I$. If $y\in I$, then $nb(I\cup\{u\})\geq nb(I\setminus\{u\})$ regardless of whether $x\in I$ In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
@@
When $w(u)<0$ and $w(u)<p(u,x)+p(u,y)$, the three vertices can not be included in $I$. $u$ can appear in $I$ iff $y$ is included and $x$ is not included in $I$. In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$ and $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
@@
Case 3: $w(u)<p(u,y)$. When $w(u)\geq 0$. $u$ can appear in $I$ iff neither $x$ nor $y$ is included in $I$. In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$ and $y\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
@@
When $w(u)<0$ and $w(u)<p(u,x)+p(u,y)$. $u$ cannot appear in $I$ because otherwise $nb(I\setminus\{u\})>nb(I)$ a contradiction. Therefore, we can remove $u$ directly.
@@
When $w(u)<0$ and $w(u)\geq p(u,x)+p(u,y)$. $u$ can appear in $I$ iff both $x$ and $y$ are included in $I$. In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\in S^{\prime}$ and $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$. ∎
@@
Observe that there exists a vertex set $I\in MGIS(G)$ that includes at least one of $x,u$ must be included in $I$. Because if none of them is included in some $I\in MGIS(G)$, then $nb(I\cup\{u\})-nb(I)\geq w(u)-\sum_{v\in N_{r}(u)}\max(0,p(u,v))\geq 0$ and therefore $I\cup\{u\}\in MGIS(G)$.
Given an arbitrary $I\in MGIS(G)$. Furthermore, $x$ and $u$ cannot be both included in $I$ since $p(u,x)=M$. In summary, there are two possible states in $I$: either $x\in I,u\notin I$ or $u\in I,x\notin I$. Therefore, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,u$ that could appear in a set in $MGIS(G)$. ∎

### v2

@@ A Technical Proofs
When $w(u)\geq p(\{u,v\})$ and $w(u)\geq 0$. If $u\notin I$ and $v\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)-p(\{u,v\})\geq nb(I)$. If $u\notin I$ and $v\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $G[I\cup\{u\}]$ does not contain permanent edges, and $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$. In summary, $u$ is included in at least one set in $MGIS(G)$.
When $w(u)<0$ and $w(u)\geq p(\{u,v\})$. If $v\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)-p(\{u,v\})\geq nb(I)$ for $u\notin I$, and $G[I\cup\{u\}]$ does not contain permanent edges. If $v\notin I$, then $nb(I\cup\{u\})<nb(I\setminus\{u\})$. In other words, there exists a set $S\in MGIS(G)$ such that $u\in S$ if and only if there exists a set $S^{\prime}\in MGIS(G^{\prime})$ such that $v\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,v$ that could appear in a set in $MGIS(G)$.
When $w(u)\geq 0$, $w(u)<p(\{u,v\})$ and $w(u)\geq\tilde{w}(v)$. If $u\notin I,v\in I$, then $nb(I\cup\{u\})\setminus\{v\})-nb(I)=w(u)-w(v)-\sum_{x\in N_{r}(v)\cap I}max(0,-p(\{v,x\}))=w(u)-\tilde{w}(v)\geq 0$, implying $I\cup\{u\}\setminus\{v\}\in MGIS(G)$. Otherwise, if $u\notin I,v\notin I$, then $nb(I\cup\{u\})-nb(I)=w(u)\geq 0$, thus $I\cup\{u\}\in MGIS(G)$. In summary, $u$ is included in at least one vertex set $I^{\prime}\in MGIS(G)$ and $v$ is not in $I^{\prime}$.
When $w(u)\geq 0$, $w(u)<p(\{u,v\})$ and $w(u)<\tilde{w}(v)$. $u$ and $v$ cannot be both included in $I$. Indeed, if $u,v\in I$, then $nb(I\setminus\{u\})=nb(I)+p(\{u,v\})-w(u)>nb(I)$. Moreover, $G[I\setminus\{u\}]$ does not contain permanent edges. This contradicts the optimality of $I$. Furthermore, if $v\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I\setminus\{u\})=w(u)\geq 0$. We conclude that there exists a vertex set in $MGIS(G)$ that includes $u$ iff there exists a vertex set in $MGIS(G^{\prime})$ that does not contain $v$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,v$ that could appear in a set in $MGIS(G)$.
Finally, when $w(u)<0$ and $w(u)<p(\{u,v\})$. We have $nb(I\cup\{u\})<nb(I\setminus\{u\})$, therefore $u$ is not in any set in $MGIS(G)$ and we can directly remove $u$. ∎
@@
Given an arbitrary $I\in MGIS(G)$. Because that $\{x,y\}\in E_{p}$, it holds that $x$ and $y$ cannot be both included in $I$.
Case 1: $w(u)\geq\max(p(\{u,x\}),p(\{u,y\}))$.
When $w(u)\geq 0$. If both $x$ and $y$ are not included in $I$ and $u\notin I$, then $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$. If $u\notin I$ and either $x\in I$ or $y\in I$, $nb(I\cup\{u\})\geq nb(I)+w(u)-\max(p(\{u,x\}),p(\{u,y\}))\geq nb(I)$. In summary, there exists a vertex set in $MGIS(G)$ that includes $u$.
When $w(u)<0$. If both $x$ and $y$ are not included in $I$, then $u\notin I$ because $nb(I\cup\{u\})=nb(I\setminus\{u\})+w(u)<nb(I\setminus\{u\})$. If either $x\in I$ or $y\in I$ and $u\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})\geq nb(I)+w(u)-\max(p(\{u,x\}),p(\{u,y\}))\geq nb(I)$. We conclude that there exists a vertex set in $MGIS(G)$ that includes $u$ iff there exists a vertex set in $MGIS(G^{\prime})$ that includes either $x$ or $y$. In other words, we can defer the decision of $u$ after determining $x$ and $y$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
Case 2: $p(\{u,y\})\leq w(u)<p(\{u,x\})$. In this case, if $y$ is in $I$, then $x\notin I$ and $I\cup\{u\}\in MGIS(G)$.
@@
When $w(u)\geq 0$ and $w(u)<\tilde{w}(x)$. There are five cases: $x\in I$ and $u,y\notin I$, $y\in I$ and $u,x\notin I$, $u\in I$ and $x,y\notin I$, $u,y\in I$ and $x\notin I$, and $u,x\in I$ and $y\notin I$. But $u,x$ cannot appear in the same $I$, otherwise $nb(I\setminus\{u\})=nb(I)+p(\{u,x\})-w(u)>0$ without containing permanent edges, which contradicts the optimality of $I$. If $y\in I$ and $u,x\notin I$, then $I\cup\{u\}\in MGIS(G)$. In summary, there exists a vertex set $I^{\prime}\in MGIS(G)$ such that either $x\in I^{\prime}$, or both $u,y\in I^{\prime}$, or $u\in I^{\prime}$. Since $u$ only connects to $x$ and $y$ in $G$, we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include $x$. In other words, we can defer $u$ after the decision of $x$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
When $w(u)<0$. If $y\notin I$, then $u\notin I$ because otherwise $nb(I\setminus\{u\})>nb(I)$. Since $u$ only connects to $x$ and $y$ in $G$, we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that includes $y$. In other words, we can defer $u$ after the decision of $y$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
Case 3: $w(u)<p(\{u,x\})$ and $w(u)<p(\{u,y\})$. When $w(u)\geq 0$, the three vertices form a triangle where every edge has penalty $M$. The result follows from theorem 3.3 of Gu et al. (2021). In particular, (1) if $w(u)\geq 0$ and $w(u)\geq\tilde{w}(x)$, then $u$ must be in at least one set in $MGIS(G)$. (2) If $w(u)\geq 0$ and $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$, then $y$ must not be in at least one set in $MGIS(G)$. There exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include $x$. (3) If $w(u)\geq 0$ and $w(u)<\tilde{w}(y)$, then there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include both $x$ and $y$. (4) When $w(u)<0$, the result follows from R5. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all conditions and all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$. ∎
@@
Case 1: $w(u)\geq p(\{u,x\})$. In this case, if one of $x,y$ is in $I$, say $x$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I)\geq w(u)-p(\{u,x\})\geq 0$ when $u\notin I$.
@@
When $w(u)\geq 0$ and $w(u)\geq p(\{u,x\})+p(\{u,y\})$. If either $x\in I$ or $y\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I)\geq w(u)-\max(p(\{u,x\})+p(\{u,y\}),p(\{u,x\}))\geq 0$ for $u\notin I$. If none of $x,y$ is in $I$, then $I\cup\{u\}\in MGIS(G)$ because $w(u)\geq 0$. Therefore, $u$ must belong to at least one set in $MGIS(G)$.
@@
When $w(u)\geq 0$ and $w(u)<p(\{u,x\})+p(\{u,y\})$. If both $x,y$ are in $I$, then $nb(I\cup\{u\})<nb(I)$. If either $x\notin I$ or $y\notin I$, then $nb(I\cup\{u\})\geq nb(I)$. In other words, including $u$ has negative profit only if both $x$ and $y$ are included in the same set in $MGIS(G)$. we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that either $x\notin S^{\prime}$ or $y\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
@@
When $w(u)<0$, it holds that $w(u)\geq p(\{u,x\})+p(\{u,y\})$ because $w(u)\geq p(\{u,x\})$ and $p(\{u,y\})<0$. If either $x\in I$ or $y\in I$, then $nb(I\cup\{u\})\geq nb(I)$. If both $x\notin I$ and $y\notin I$, then $nb(I\cup\{u\})<nb(I)$. In other words, including $u$ has positive profit if and only if either $x$ or $y$ is included. we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that either $x\in S^{\prime}$ or $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
Case 2: $p(\{u,y\})\leq w(u)<p(\{u,x\})$. In this case, we consider whether the three vertices $u,x,y$ can be included in $I$.
@@
When $w(u)\geq 0$ and $w(u)\geq p(\{u,x\})+p(\{u,y\})$, the all three vertices may be included in $I$. All combinations of $x,y,u$ may appear in $I$. Except that if $x$ is included and $y$ is not included in $I$, then $u\notin I$ because $nb(I\cup\{u\})=nb(I\setminus\{u\})+w(u)-p(\{u,x\})<nb(I\setminus\{u\})$. In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that either $x\notin S^{\prime}$ or $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
@@
When $w(u)\geq 0$ and $w(u)<p(\{u,x\})+p(\{u,y\})$, the three vertices can not be included in $I$. Moreover, $u$ and $x$ cannot be both included by $I$ because otherwise $nb(I\setminus u)>nb(I)$ without containing permanent edges, which contradicts the optimality of $I$. If $x\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})\geq nb(I\setminus\{u\})$. In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
@@
When $w(u)<0$ and $w(u)\geq p(\{u,x\})+p(\{u,y\})$, the three vertices can be included in $I$. If $y\notin I$, then $nb(I\cup\{u\})\leq nb(I\setminus\{u\})$ regardless of whether $x\in I$. If $y\in I$, then $nb(I\cup\{u\})\geq nb(I\setminus\{u\})$ regardless of whether $x\in I$ In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
@@
When $w(u)<0$ and $w(u)<p(\{u,x\})+p(\{u,y\})$, the three vertices can not be included in $I$. $u$ is in $I$ iff $y$ is included and $x$ is not included in $I$. In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\notin S^{\prime}$ and $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
@@
Case 3: $w(u)<p(\{u,y\})$. When $w(u)\geq 0$. $u$ is in at least one set $I\in MGIS(G)$ iff neither $x$ nor $y$ is included in $I$. In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\notin S^{\prime}$ and $y\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
@@
When $w(u)<0$ and $w(u)<p(\{u,x\})+p(\{u,y\})$. $u$ is not in any $I\in MGIS(G)$ because otherwise $nb(I\setminus\{u\})>nb(I)$ without containing permanent edges, a contradiction. Therefore, we can remove $u$ directly.
@@
When $w(u)<0$ and $w(u)\geq p(\{u,x\})+p(\{u,y\})$. $u$ is in at least one set $I\in MGIS(G)$ iff both $x$ and $y$ are included in $I$. In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\in S^{\prime}$ and $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$. ∎
@@
Observe that there exists a vertex set $I\in MGIS(G)$ that includes at least one of $x,u$ must be included in $I$. Because if none of them is included in at least one set $I\in MGIS(G)$, then $nb(I\cup\{u\})-nb(I)\geq w(u)-\sum_{v\in N_{r}(u)}\max(0,p(\{u,v\}))\geq 0$ and therefore $I\cup\{u\}\in MGIS(G)$.
Given an arbitrary $I\in MGIS(G)$. Furthermore, $x$ and $u$ cannot be both included in $I$ since $\{u,x\}\in E_{p}$. In summary, there are two possible states in $I$: either $x\in I,u\notin I$ or $u\in I,x\notin I$. Therefore, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,u$ that could appear in a set in $MGIS(G)$. ∎